# Offload Rethinking by Cloud Assistance for Efficient Environmental Sound Recognition on LPWANs

### v1

@@ Abstract.
Embedded intelligence, cloud offloading, environmental sound classification, batteryless computing, LPWANs
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organizationccs: Embedded and cyber-physical systemsccs: Embedded systems
@@ 1. Introduction
Despite their potential, edge devices’ limited processing power and energy constraints make high-accuracy on-device predictions challenging (lee2019intermittent, ; gobieski2019intelligence, ; jeon2023harvnet, ; islam2020zygarde, ). To overcome these limitations, recent efforts focus on edge-cloud collaboration (ahn2024split, ; huang2023rethink, ; yao2020deep, ; hojjat2024limitnet, ). Compared to on-device learning, edge-cloud collaboration offloads inference tasks to the cloud server via wireless networks like Bluetooth LE, Wi-Fi, and LTE (ahn2024split, ; huang2023rethink, ; yao2020deep, ; hojjat2024limitnet, ). This improves accuracy and reduces latency, but adds network dependency and energy overheads of costly wireless communication. Additionally, network coverage is a key factor in matching the diverse application scenarios. LTE provides wide coverage but is power-hungry for resource-constrained devices. In contrast, BLE and Wi-Fi are energy-efficient but have limited range, making them unsuitable for wide-area sensing applications. For instance, biologists deploy sensor nodes for acoustic bird species identification across vast rainforest areas (verma2024aviear, ). Similarly, civil engineers and researchers use sensor nodes for large-scale acoustic sensing and classification in urban noise monitoring (tan2021extracting, ). Although not sufficiently explored in previous studies, we find Low-Power Wide-Area Networks (LPWANs) ideal for these scenarios, balancing range and communication efficiency in edge-cloud collaborations.
@@
In this paper, we introduce a novel resource-aware cloud-assisted machine learning system, ORCA (Offload Rethinking by Cloud Assistance), tailored for environmental sound recognition on ultra-low-power batteryless devices operating over LPWANs like LoRa networks. First, we leverage LPWANs to enable low-power, wide-area, and long-range edge-cloud collaboration for wide-area sound sensing applications. Second, we propose a novel cloud assistance strategy for high-accuracy, resource-efficient, and LPWAN-adapted cloud offloading. In cloud assistance, instead of on-device inference as Figure 1(a) or offloading by splitting the pipeline across wireless channels and generating inference results on the cloud as Figure 1(b), ORCA keeps the entire inference pipeline on edge devices while leveraging the server for assistance. Specifically, the server identifies the most important input features sampled and uploaded by the edge devices. Here, we focus on identifying the most important frequency-domain features as previous studies show their effectiveness in environmental sound classification (phaye2019subspectralnet, ). After that, the server provides the edge device with feature importance information through the downlink. This process is shown as the red arrows in Figure 1(c). The edge devices can now use the given information to selectively process and perform inference on the most important features as the blue arrow in Figure 1(c), thereby saving energy and time by scaling down the inputs and improving accuracy. As a result, the role of the server is now ”assisting” edge devices for inference, avoiding edge devices being dependent on it. Specifically, ORCA tackles three key challenges of cloud assistance over LPWANs:
@@
(1) High communication costs and low data rate: Though LPWANs like LoRa are “low-power”, wireless communication remains resource-intensive, e.g., 40$\times$ for uplink and 6$\times$ for downlink to the average power of on-device inference (mileiko2023run, ). Additionally, LPWANs operate at a low data rate, e.g., 0.3-5.5 kbps in 125 kHz LoRa channel (bor2016lora, ). Therefore, previous data streaming on BLE, Wi-Fi, or LTE (huang2023rethink, ; yao2020deep, ) is infeasible. In ORCA, we tackle this problem with cloud assistance, transmitting a low-resolution spectrogram from edge to cloud and selecting important features. ORCA implements this process by feeding the low-resolution spectrogram into a pre-trained vision transformer and extracting the feature importance masks from the self-attention maps. Our results show that the low-resolution spectrogram not only saves payloads but is also effective in learning the importance of features.
@@
Contributions: (1) We propose a novel cloud assistance strategy for environmental sound recognition on resource-limited embedded systems and LPWANs. By leveraging vision transformer architecture for subspectral feature selection and efficient on-device spectral encoding convolutional neural networks (CNNs), ORCA achieves 2.5–12.5 p.p. accuracy improvements on five public environmental sound datasets. (2) We provide optimized assistance resolution strategies under dynamic communication costs and packet losses to ensure resource-efficient cloud assistance in dynamic environments. (3) We build a real-world testbed to evaluate accuracy, payload size, energy consumption, inference latency, and system overhead using collected LoRa traces. ORCA achieves up to 80$\times$ energy savings and 220$\times$ latency reduction, while maintaining accuracy comparable to state-of-the-art methods. We release an open-source LoRa library for batteryless computing on the ultra-low-power MSP430 microcontroller [^1] Library available at: https://anonymous.4open.science/r/508-SenSys25-ORCA/README.md.
2. Background and Related Works
@@
ORCA Cloud assistance LoRa Low-Res specs Low resolution 0-0.1KB
@@ 3. Preliminary Study
Figure 4. ORCA cloud-assisted design overview.
@@ 4.4. Cloud-Assisted Inference
The next challenge is to create a multi-resolution representation for inference. As discussed in preliminary studies in Section 3, discriminative information varies between spectral bands of the spectrogram. With the full low-resolution spectrogram available from preprocessing, we use two 2-layer shallow CNN as encoder encoders, one for low resolution and one for high resolution. The encoded features are fused channel-wise into a single vector and fed into the Multi-Res classifier for final classification. This architecture reduces inference costs by leveraging spectrograms at different resolutions. If cloud assistance is unavailable, an additional Single-Res classifier is employed to process the output of the Low-Res encoder only. All components are pre-trained offline in a two-stage supervised process. First, we train the low-res encoder, high-res encoder, and multi-resolution classifier together with the attention masks generated by the pre-trained cloud vision transformer. In the second stage, we freeze all other components and train the single-resolution classifier independently.
@@ 4.5. Resource-Aware Scheduler
Figure 8. Cloud-assisted offloading flow chart.
@@ 6.1. Testbed Implementation
Software: The PyWavelets library (Lee2019PyWavelet, ) is employed for preprocessing the audio clips, while PyTorch (paszke2019pytorch, ) is used for pretraining the server and edge models. Raspberry Pi server also uses PyTorch (paszke2019pytorch, ) for cloud assistance. Additionally, we have developed a custom wavelet transform library and an inference engine integrated with the Low-Energy Accelerator (LEA) on the MSP430 (MSPLEA2016, ), aimed at accelerating both preprocessing and inference processes. For the server side, we use the Arduino RadioHead library (mccauley2013radiohead, ). For MSP430, We adapt the RadioHead library (mccauley2013radiohead, ) over the SPI interface to facilitate communication with the RFM95W transceiver for both uplink and downlink. To the best of our knowledge, this is the first usable and open-source implementation of the LoRa library for the MSP430.
@@ 6.3. Payload Size
Table 2. On-device energy measurement and end-to-end inference latency for preprocessing, downlink, server processing, and inference.
@@ 8. Conclusion
Acknowledgements
This work was supported in part by the National Science Foundation under Grants #2112665 (TILOS AI Research Institute), #2003279, #1911095, #1826967, #2100237, #2112167, and in part by PRISM and CoCoSys, centers in JUMP 2.0, an SRC program sponsored by DARPA.

### v2

@@ Abstract.
Embedded intelligence, cloud offloading, environmental sound classification, batteryless computing, LoRa, LPWANs
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 6–9, 2025; Irvine, CA, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Sensor networksccs: Computer systems organization Embedded systemsccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Despite their potential, edge devices’ limited processing power and energy constraints make high-accuracy on-device predictions challenging (lee2019intermittent, ; gobieski2019intelligence, ; jeon2023harvnet, ; islam2020zygarde, ). To overcome these limitations, recent efforts focus on edge-cloud collaboration (ahn2024split, ; huang2023rethink, ; yao2020deep, ; hojjat2024limitnet, ). Compared to on-device learning, edge-cloud collaboration offloads inference tasks to the cloud server via wireless networks like Bluetooth LE, Wi-Fi, and LTE (ahn2024split, ; huang2023rethink, ; yao2020deep, ; hojjat2024limitnet, ). This improves accuracy and reduces latency, but adds network dependency and energy overheads of costly wireless communication. Additionally, network coverage is a key factor in matching the diverse application scenarios. LTE provides wide coverage but is power-hungry for resource-constrained devices. In contrast, BLE and Wi-Fi are energy-efficient but have limited range, making them unsuitable for wide-area sensing applications. For instance, biologists deploy sensor nodes for acoustic bird species identification across vast rainforest areas (verma2024aviear, ). Similarly, civil engineers and researchers use sensor nodes for large-scale acoustic sensing and classification in urban noise monitoring (tan2021extracting, ). We find Low-Power Wide-Area Networks (LPWANs) ideal for these scenarios, balancing range and communication efficiency in edge-cloud collaborations.
@@
In this paper, we introduce a novel resource-aware cloud-assisted machine learning system, ORCA (Offload Rethinking by Cloud Assistance), tailored for environmental sound recognition on ultra-low-power batteryless devices operating over LPWANs like LoRa networks. First, we leverage LPWANs to enable low-power, wide-area, and long-range edge-cloud collaboration for wide-area sound sensing applications. Second, we propose a novel cloud assistance strategy for high-accuracy, resource-efficient, and LPWAN-adapted cloud offloading. In cloud assistance, instead of on-device inference as Figure 1(a) or offloading by splitting the pipeline across wireless channels and generating inference results on the cloud as Figure 1(b), ORCA keeps the entire inference pipeline on edge devices while leveraging the server for assistance. Specifically, the server selects the most important features uploaded by the edge devices. Here, we focus on identifying the most important frequency-domain features as previous studies show their effectiveness in environmental sound classification (phaye2019subspectralnet, ). After that, the server provides the edge device with feature importance information through the downlink. This process is shown as the red arrows in Figure 1(c). The edge devices can now use the given information to selectively process and perform inference on the most important features as the blue arrow in Figure 1(c), thereby saving energy and time by scaling down the inputs and improving accuracy. As a result, the role of the server is now ”assisting” edge devices for inference, avoiding edge devices being dependent on it. Specifically, ORCA tackles three key challenges of edge-cloud collaboration over LPWANs:
@@
(1) High communication costs and low data rate: Though LPWANs like LoRa are “low-power”, wireless communication remains resource-intensive, e.g., 40$\times$ for uplink and 6$\times$ for downlink to the average power of on-device inference (mileiko2023run, ). Additionally, LPWANs operate at a low data rate, e.g., 0.3-5.5 kbps in 125 kHz LoRa channel (bor2016lora, ). As a result, previous data streaming approaches on BLE, Wi-Fi, or LTE (huang2023rethink, ; yao2020deep, ) are infeasible. To further reduce payload sizes, we propose transmitting low-resolution spectrograms for cloud processing. However, our preliminary studies indicate that while low-resolution spectrograms significantly reduce payloads compared to conventional cloud offloading methods, they also degrade cloud inference accuracy. On the other hand, we observe that feature selection is more resilient to accuracy degradation under low-resolution and limited communication budgets compared to cloud inference. Therefore, we introduce a cloud-assisted approach that transmits low-resolution spectrograms to the cloud for feature selection rather than inference. ORCA implements this by feeding the low-resolution spectrogram into a pre-trained vision transformer and extracting feature importance masks from the self-attention maps.
@@
Contributions: (1) We propose a novel cloud assistance strategy for environmental sound recognition on resource-limited embedded systems and LPWANs. By leveraging vision transformer architecture for subspectral feature selection and efficient on-device spectral encoding convolutional neural networks (CNNs), ORCA achieves 2.5–12.5 p.p. accuracy improvements on five public environmental sound datasets.
(2) We provide optimized assistance resolution strategies under dynamic communication costs and packet losses to ensure resource-efficient cloud assistance in dynamic environments.
(3) We build a real-world testbed to evaluate accuracy, payload size, energy consumption, inference latency, and system overhead using collected LoRa traces. ORCA achieves up to 80$\times$ energy savings and 220$\times$ latency reduction, while maintaining accuracy comparable to the state-of-the-art methods. We release an open-source LoRa library for batteryless computing on the ultra-low-power MSP430 microcontroller [^1] Library available at: https://github.com/lezhangleonard/ORCA.
@@
ORCA Cloud assistance LoRa Spectrogram Low resolution 0-0.1KB
@@ 4. ORCA Design
Figure 4. ORCA cloud-assisted design overview.
@@ 4.4. Cloud-Assisted Inference
The next challenge is to create a multi-resolution representation for inference. As discussed in preliminary studies in Section 3, discriminative information varies between spectral bands of the spectrogram. With the full low-resolution spectrogram available from preprocessing, we use two 2-layer shallow CNN as encoders, one for low resolution and one for high resolution. The encoded features are fused channel-wise into a single vector and fed into the Multi-Res classifier for final classification. This architecture reduces inference costs by leveraging spectrograms at different resolutions. If cloud assistance is unavailable, an additional Single-Res classifier is employed to process the output of the Low-Res encoder only. All components are pre-trained offline in a two-stage supervised process. First, we train the low-res encoder, high-res encoder, and multi-resolution classifier together with the attention masks generated by the pre-trained cloud vision transformer. In the second stage, we freeze all other components and train the single-resolution classifier independently.
@@ 4.5. Resource-Aware Scheduler
Figure 8. Cloud-assisted offloading flow chart.
@@ 6.1. Testbed Implementation
Software: The PyWavelets library (Lee2019PyWavelet, ) is employed for preprocessing the audio clips, while PyTorch (paszke2019pytorch, ) is used for pretraining the server and edge models. Raspberry Pi server also uses PyTorch (paszke2019pytorch, ) for cloud assistance. Additionally, we have developed a custom wavelet transform library and an inference engine (zhang2022demo, ) integrated with the Low-Energy Accelerator (LEA) on the MSP430 (MSPLEA2016, ), aimed at accelerating both preprocessing and inference processes. For the server side, we use the Arduino RadioHead library (mccauley2013radiohead, ). For MSP430, We adapt the RadioHead library (mccauley2013radiohead, ; eccob2020msp430, ) over the SPI interface to facilitate communication with the RFM95W transceiver for both uplink and downlink. To the best of our knowledge, this is the first usable and open-source implementation of the LoRa library for the MSP430.
@@ 6.3. Payload Size
Table 2. On-device energy measurement and end-to-end inference latency for each stage. Uplink is a variable for optimization and mentioned in the text.
@@ 8. Conclusion
Acknowledgements.
This work was supported in part by the National Science Foundation under Grants #2112665 (TILOS AI Research Institute), #2003279, #1911095, #1826967, #2100237, #2112167, and in part by PRISM and CoCoSys, centers in JUMP 2.0, an SRC program sponsored by DARPA. We thank the reviewers and the shepherd for their insightful comments and suggestions.